OpenStudy (anonymous):

PLZ HELP I JUST NEED HELP NOT THE ANSWER PLZ ANYONE simplify (6x^-2)^2(0.5x)^4

OpenStudy (anonymous):

ill award anyone who FULLY HELPS ME @StudyGurl14 @sparrow2 @Study312 @skullpatrol

OpenStudy (studygurl14):

We already did this one, remember?

OpenStudy (anonymous):

was the answer 2.25?

OpenStudy (studygurl14):

Yes

OpenStudy (study312):

@StudyGurl14 I got a question for u. What grade is this kind of math?

OpenStudy (anonymous):

1. 6​x^-2*6x^-2=36x^-4 2. (0.5x)^4=(0.5x)*(0.5x)*(0.5x)*(0.5x)=0.0625x 3 (0.0625x(36x^-4)=2.25 is this right?

OpenStudy (studygurl14):

No specific grade. It's algebra though

OpenStudy (studygurl14):

The second step is incorrect, and so is the 4th

OpenStudy (studygurl14):

the 3rd I meant

OpenStudy (studygurl14):

\(\Large (0.5x)^4 = 0.5^4x^4=0.0625x^4\)

OpenStudy (anonymous):

how two fours though? just asking so I can understand

OpenStudy (anonymous):

@StudyGurl14